A 90.0-gram lead bullet hits a steel plate, both initially at 29 °C. The bullet melts and splatters on impact. Assuming that the bullet receives 76% of its kinetic energy as heat energy, at what minimum speed must it be traveling to melt on impact? (The melting temperature of Lead is 327°C)
Answer #1

Heat, Q = m C ΔT + mL

Q = (0.09 x 130 x 298) + (0.09 x 22400)

Q = 5502.6 J

KE = 5502.6/0.76 = 7240.26 J

0.5 x 0.09 v^2 = 7240.26

Velocity, v = 401.12 m/s

Comment in case any doubt please rate my answer....